OpenStudy (anonymous):

Solve the system by the substitution method: 2x+8y=3 x=8-4y

OpenStudy (anonymous):

you want answer or solution?

OpenStudy (anonymous):

Solution please

OpenStudy (anonymous):

Substitute your second equation in first..i mean like this... 2x + 8y = 3 2(8-4y) + 8y = 3 16 - 8y = +8y = 3 16 = 3 what??? okay, the lines are parallel. there is no solution..

OpenStudy (anonymous):

the third line of solving is 16 - 8y +8y = 3

OpenStudy (anonymous):

umm..did you understand?

OpenStudy (anonymous):

so y=13?

OpenStudy (anonymous):

and x=-44?

OpenStudy (anonymous):

No..there is no intersection point. They are parallel lines. They don't intersect.

OpenStudy (anonymous):

yeah i got that! thanks!